You are given data for a hypothetical simple closed economy;



C= 250+0.85Y



I= 150



G=200



Where C is consumer expenditure, Y is national income, G is government expenditure on goods and services and I is investment expenditure. All amounts are in K million.



1. Derive a savings function using the information for the hypothetical economy

1
Expert's answer
2021-12-17T18:17:11-0500

S=Y-C

Y=C+I+G=250+0.85Y+150+200

0.15Y=600

Y=4000

"C=250+0.85\times4000=3650"

S=4000-3650=350